Carbon monoxide and oxygen react in the following manner:


2 CO (g) + O2 (g) → 2 CO2 (g)

Information on the reaction is given in the table below:

[CO] 0.020 M, 0.040 M , 0.020 M , 0.011 M , 0.59 M , ♥

[O2] 0.020 M , 0.020 M, 0.040 M , 1.22 M , ♣ , 0.59 M

Rate 3.68 x 10-5 M/min , 1.47 x 10-4 M/min, 7.36 x 10-5 M/min , ♠ , 8.62 x 10-3 M/min , 8.62 x 10-3 M/min


Part A

Determine the order of the reaction with respect to CO.

Part B

Determine the order of the reaction with respect to O2.

Part C

Calculate "k".

Part D

Calculate ♠.

Part E

Calculate ♣.

Part F

Calculate ♥.


1
Expert's answer
2021-04-25T23:11:17-0400

Part A

Rate="8.9\times 10^4[0.020M]^3[0.040M]"

"=0.0285M\/min"

Part B

"=8.9\times 10^4[0.020M]^3[1.44M]"

"=1.025M\/min"

Part C

"k=0.0285" "M\over min" "=k(0.020M)^3(0.040M)"

"k=" "{{0.0285}M\over min}\over (0.020M)^3 (0.040M)"

"=8.9\times 10^4" "M^3\over min"

Part D

"1\over 4" "\times 0.0285M/min" "=7.125\times 10^{-3} M/min"

Part E

"O_2=" "0.020M\over 1.4min" "=0.0143M\/min"

Part D

"CO_2=" "0.040M\over 1.4min" "=0.0285M\/min"